# Optimal Convex Cover as Collision-free Space Approximation for Trajectory Generation

### v2

@@ IV-A Initialization and Polytopes Generation
Given an initial path $p^{0}$, we upsample waypoints by adding auxiliary intermediate waypoints between consecutive original waypoints to avoid creating excessively long ellipsoids that exceed a specified distance threshold $\alpha$. We use the procedure InitEllipsoids that takes as input line segments of the initial path and outputs the ellipsoid with $d_{i}=(p_{i}+p_{i-1})/2$ and
$L_{i}L_{i}^{T}=v_{i}v_{i}^{T}\frac{||p_{i}-p_{i-1}||_{2}^{2}}{4}+(I-v_{i}v_{i}^{T})\epsilon^{2}.$ (10)
@@
${\rm s.t.}$ $\mathcal{E}_{i}\subset\mathcal{P}_{i},\mathcal{P}_{i}\cap\mathcal{O}=\varnothing\quad\forall i=1,\cdots,M,$ (11b)
We use a similar method in [3] to solve this problem by finding tangent separating planes from the ellipsoid. To save computation, we define a local bounding box by adding a margin of ±$\pm l$ around the ellipsoid for each polytope generation. This margin ensures that only the local obstacles within this expanded range are considered. The overlap condition $\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\neq\varnothing$ is satisfied by construction. Additionally, we store the obstacle points near each initialized ellipsoid for parallel updating of polytopes during iteration.
@@ IV-B Waypoints Update with Trajectory Heuristics
IV-B1 Minimum Path Length.
@@
where constraint (13c) ensures that waypoints are generated inside the overlapping region of every two consecutive polytopes.
@@
IV-B2 Minimum Control Effort.
@@ IV-C Maximum Volume Ellipsoids with Inner&Outer Constraints
Figure 4: Demonstration of geometric constraints for an ellipsoid (blue). (a) The general formulation of inner and outer Löwner-John ellipsoid. (b) The Inner Löwner-John ellipsoid with outer line segment constraints. (c) The Inner Löwner-John ellipsoid with outer convex bodies constraints. The dash shapes represent neighbored polytopes.
Given current waypoints $p^{k}$, we can formulate the sub-problem of finding the $i$-th ellipsoid by
@@ V-C Comparison of Convex Cover
Figure 9: The illustration of SFC generated by three methods. More comparisons are available at https://www.youtube.com/watch?v=J3YmIk_TKiQ
To further evaluate the performance, we conduct 10,000 additional tests in a broader set of simulated environments with dimensions of $30\times 30\times 5$m3 characterized by the ECS (density index, cluster index, and structure index). The results are filtered to exclude all invalid cases encountered by any of the methods. We categorize test cases based on the range of the density index to represent environment complexity and show the results in Fig. 10.
Figure 10: Comparisons with increasing density index (defined in [26]).
As the density index increases, environments become more complex, which leads to reduced polytope volumes and decreased trajectory performance across all three methods. When the density index exceeds 0.5, finding a feasible geometric path over long distances becomes challenging, resulting in shorter paths and fewer segments.
Figure 11: Comparisons with increasing start-end distance.
Benchmark results classified by start-end distance (the Euclidean distance between the start to end point) are demonstrated in Fig. 11. As the start-to-end distance increases, our method demonstrates improved performance in terms of computation time, trajectory cost, and path length. It is clear that the performance of any algorithm depends on the complexity of the environment and start-to-end distance, and it is difficult to draw conclusions based on average performance. However, as shown in Fig. 10 and 11, the worst-case cost for our proposed method is lower than that obtained by RILS and IRIS with similar computational time. More importantly, this is the first systematic analysis of multiple factors that influence the trajectory cost: the waypoint and number of segments in the front-end plan, the volumes of the polytopes in the convex cover, and the volumes of consecutive polytopes.
@@ VI Conclusion and Future Work
This paper introduces a novel algorithm to find an optimal convex cover of collision-free space as SFC for trajectory optimization. The convex cover is formulated as a joint optimization of ellipsoids, polytopes, and intermediate waypoints, which are updated iteratively to find optimal solutions. The proposed method can be applied to two-stage motion planning as an intermediate process, or optimizing with trajectory generation jointly. Our future work will address whole-body trajectory planning for $SE$(2) and $SE$(3).

### v3

@@ IV-A Initialization and Polytopes Generation
Given an initial path $p^{0}$, we upsample waypoints by adding auxiliary intermediate waypoints between consecutive original waypoints to avoid creating excessively long ellipsoids that exceed a specified distance threshold $\alpha$. We use the procedure InitEllipsoids that takes as input line segments of the initial path and outputs the ellipsoid centered at $d_{i}=(p_{i}+p_{i-1})/2$. Let $\mu_{i}=\frac{p_{i}-p_{i-1}}{||p_{i}-p_{i-1}||{2}}$ be the unit vector along the displacement from waypoint $p_{i-1}$ to waypoint $p_{i}$. Then $L_{i}$ is chosen so that
$L_{i}L_{i}^{T}=\mu_{i}\mu_{i}^{T}\frac{||p_{i}-p_{i-1}||_{2}^{2}}{4}+(I-\mu_{i}\mu_{i}^{T})\epsilon^{2}.$ (9)
@@
${\rm s.t.}$ $\mathcal{E}_{i}\subset\mathcal{P}_{i},\mathcal{P}_{i}\cap\mathcal{O}=\varnothing\quad\forall i=1,\cdots,M.$ (10b)
We adopt a method similar to [4] to solve this problem by finding tangent separating planes from the ellipsoid. To save computation, we define a local bounding box by adding a margin of ±$\pm l$ around the ellipsoid for each polytope generation. This margin ensures that only the local obstacles within this expanded range are considered. The overlap condition $\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\neq\varnothing$ is satisfied by construction. Additionally, we store the obstacle points near each initialized ellipsoid for parallel updating of polytopes during iteration.
@@ IV-B Waypoints Update with Trajectory Heuristics
IV-B1 Minimum Path Length
@@
where constraint (12c) ensures waypoints are generated within the overlapping region of every two consecutive polytopes.
@@
IV-B2 Minimum Control Effort
@@ IV-C Maximum Volume Ellipsoids with Inner&Outer Constraints
Figure 4: Demonstration of geometric constraints for an ellipsoid (blue). (a) The general formulation of inner and outer Löwner-John ellipsoid. (b) The Inner Löwner-John ellipsoid with outer line segment constraints.
Given current waypoints $p^{k}$, we can formulate the sub-problem of finding the $i$-th ellipsoid $\mathcal{E}_{i}(L_{i},d_{i})$ by
@@ V-C Comparison of Convex Cover
Figure 9: The illustration of SFC generated by three methods. More comparisons are available at https://youtu.be/k7CI7-fgSXE
It is clear that the performance of any algorithm is influenced by the complexity of both the environment and the task, which makes it difficult to draw conclusions based on average performance. To further compare with benchmark methods, we conduct 10,000 additional tests in simulated environments characterized by the ECS (density index, cluster index, and structure index). The results are filtered to exclude cases where any method fails to find feasible corridors. We categorize test cases based on the range of the density index to represent environment complexity and show the results in Fig. 10.
Figure 10: Benchmark comparison with different environmental parameters, including density index, cluster index, structure index (as defined in [27]) and start-to-end distance.
Overall, our method achieves lower trajectory costs across various environments. As density and cluster indices increase, environments become more complex, which leads to increased trajectory costs across all three methods. In contrast, more structured environments allow for trajectory generation with lower costs. The comparison results categorized by start-to-end distance (the Euclidean distance between the start and end points), show that our method outperforms others as the distance increases. More importantly, this is the first systematic analysis of multiple factors that influence the trajectory cost: the waypoint and number of segments in the front-end plan, the volumes of the polytopes in the convex cover, and the volumes of consecutive polytopes.
@@ VI Conclusion and Future Work
This paper introduces a novel algorithm for optimizing the convex cover of collision-free space as SFC for trajectory optimization. The convex cover is formulated as a joint optimization of ellipsoids, polytopes, and intermediate waypoints, which are updated iteratively to find optimal solutions. The proposed method can be applied to two-stage motion planning as an intermediate process, or optimizing with trajectory generation jointly. Our future work will address whole-body trajectory planning for $SE$(2) and $SE$(3).